If the gauge number of a wire is increased by 6, t
Answer #1

The resistance of a wire is proportional to its length and inversely proportional to the square of its diameter.

Since the wire gauge is going from 7 to 24 (difference of 17) and the diameter is halved for every increase of 6, the diameter for this change is reduced to (1/2^1/6)^17 =0.140

(1.21Ω)(1)^2 / (930 ft.) = (R)(0.140)^2 / (31 ft.)
R = 2.057Ω